OpenStudy (anonymous):

An amphitheater has 12 seats in the first row, 19 in the second row, and 26 in the third row. If the pattern continues, how many seats are in the 13th row?

OpenStudy (curry):

96

OpenStudy (anonymous):

this is an arithmetic series with common difference 7 13th term = a + 12m = 12 + 12*7 = 96

OpenStudy (anonymous):

d not m

OpenStudy (anonymous):

series {12, 19, 26} term1- terms 2= -7 term2- term 3= -7 so it's an arithmetic progression t(x) = 7( x-1) +12 the next term is the 4th term so 7(4-1)+12=33